# Jan Schiefelbein-Lach4, Oliver Pohl5, Andreas Ulbig2, and Martin Grohe1

### v1

@@ IV Experiments
To enlarge our training set, we performed data augmentation, where we added noise to the bounds for $P^{\text{min}}$, $P^{\text{max}}$. The set points of each augmented bus are set to its maximum. We augment cases with violation of the lower voltage band ten times in total to increase their representation. We kept the unaugmented supply task as test data and trained only on the augmented data. We train on shuffled data in batches of each 20 supply tasks. The model gets each time-step five times in a row. Except for the first time, the set-points were adapted according to the model’s previous decision, so the model has the chance to correct previous decisions. We trained for 1,200,000 steps, so 240,000 different training samples (of augmented data), each five times. We run our experiments on a desktop computer with an AMD Ryzen 7 7700X processor and an NVIDIA GeForce RTX 4070 Ti GPU.
For our baselines, we take the state estimation from pandapower [22] together with the OPF solver from pypower [20] and the OPF with complete information. Since our state estimation is relatively exact, there are only minor differences between se+opf and the OPF alone. Thus, we only depict the results of se+opf.
In Figure 2 are depicted the test results of our model in comparison to baseline and se+opf (state estimation + OPF). In this case, we call baseline the case where every bus with flexibility is on its maximum set point. Each of the scatter plots contains all supply tasks. On the $x$ axes, the relative curtailment of the current grid state is visualised. The relative curtailment is the ratio of the sum of absolute power output or input, respectively, at each controllable bus and the maximal flexibility. The colour of each point corresponds to maximal flexibility in kW. On the $y$ axes, we plot the upper voltage band (top), the lower voltage band (middle), and the relative asset load (bottom), each with their corresponding limits (blue dashed line). In the baseline column, there is no curtailment visible, but several violations. In the se+opf column, there has been curtailment, which fixes all the violations. Starting in the upper voltage band, the limit is at 1.05p.u. All cases are below that line, even under 1.04p.u. In the plot of the lower voltage band, there are again no violations remaining, and the lowest points are all on the 0.95p.u. line. A similar behaviour can be seen in the plot for asset overload, where many points are now lying on the 100% line. Last, the reinforcement learning model shows a similar behaviour as se+opf. However, there is much more noise in the data. Except for 22 cases in the asset overload, all violations have been resolved by the RL agent. Though the RL agent detects these 22 cases as violations, it does not curtail them sufficiently. In a direct comparison of se+opf and the RL model, there are substantial similarities: in the upper voltage band, there are clusters of points in different lines, which are visible in both plots. In the lower voltage band, clusters of points appear in similar regions, although with a more extensive spread. The same behaviour can be seen in the plot for asset overload. These clusters correspond to the different strands of the grid. The RL model shows much more noise on these “structures” and more curtailment. We also analysed our model’s output for $Q$, which is close to the output of se+opf.
@@
V Discussion
@@
VI Conclusion
We introduce a machine learning model for solving state estimation and OPF in end-to-end fashion. We train and validate our model on real grid data and show that the model is able to detect and counteract violations in most cases, while being faster than traditional solvers. Although the model does not provide a sufficient amount of curtailment in some cases, it correctly detects all violations. This drawback could be mitigated with more training data and further fine-tuning of the hyperparameters. For future work, the model could be extended to quantify the uncertainty of its output. In grid operation, the decisions of the models would only be applied if the model is certain of its decision. The overall accuracy could be improved by using ensembles of embeddings, where the outputs are ranked by their (un)certainty.

### v2

@@ III-B Training
We train on shuffled data in batches of each 20 supply tasks. The model gets each time-step five times in a row. Except for the first time, the set points were adapted according to the model’s previous decision, so the model has the chance to correct previous decisions. We trained for 1,200,000 steps, so 240,000 different training samples (of augmented data), each five times. We run our experiments on a desktop computer with an AMD Ryzen 7 7700X processor and an NVIDIA GeForce RTX 4070 Ti GPU.
III-C Evaluation
To evaluate the models, we compare the results of the RL agent (rl) with the OPF (opf). However, complete information is required for the OPF, so a state estimation is necessary before using it.
In the following figures, test results of our model in comparison to the opf are depicted. On the $x$ axes, the relative curtailment of the current grid state is visualised. The relative curtailment is the ratio of the sum of absolute power output or input, respectively, at each controllable bus and the maximal flexibility. The colour of each point corresponds to the available flexibility available in the grid in kW. Each point in the plots represents a single supply task on the grid topology and shows the worst-case situation on the grid (highest asset loading or lowest/highest voltage).
In Fig. 3, the asset overloading in the grid is presented. The opf solution presents an optimal curtailment of all flexibilities to maintain the grid within the operational limits. The curtailment is along the 100% line to solve all critical asset loadings cost optimal. The rl decision shows similar behaviour but in a suboptimal way. The agent’s decisions are, in a way, overachieving and curtailing increasingly more to stay within the permissible physical bounds but with way less available data than the OPF.
Figure 3: Asset Overloading and Relative P Curtailment
When looking at the lower voltage band in Fig. 4, there are few curtailment decisions done by the opf to stay within the voltage band of 0.95 p.u. and 1.05 p.u. The rl also stays within the voltage band, but a similar suboptimal decision can be observed.
Figure 4: Lower Voltage Band and Relative P Curtailment
Lastly, we looked at the relative Q curtailment in Fig. 5 to compare the decisions of the rl agent with the opf. It is notable that the rl agent learned to use $Q$ similarly as the opf to achieve congestion-free grid states.
Figure 5: Asset Overloading and Relative Q Curtailment
Except for 22 cases in the asset overload, all violations have been resolved by the rl agent. Though the rl agent detects these 22 cases as violations, it does not curtail them sufficiently. The rl agent shows much more noise and more curtailment but learns to curtail with just limited information with a high accuracy and high detection rate of critical grid states. When comparing the computation times II, the end-to-end learning approach outperforms the state of the methods.
@@ III-C Evaluation
Especially when comparing the rl agent with the OPF with an initial State Estimation, the computation time increases significantly.
IV Discussion
@@
V Conclusion
We introduce a machine learning model for solving state estimation and OPF in an end-to-end fashion. This grid decision learning enables the curative curtailment of flexibilities. We train and validate our model on real grid data and show that the model is able to detect and counteract violations in most cases, while being faster than traditional solvers. Although the model does not provide sufficient curtailment in some cases, it correctly detects all violations. This drawback could be mitigated with more training data and further fine-tuning of the hyperparameters.
For future work, the model could be extended to quantify the uncertainty of its output. In grid operation, the models’ decisions are only applied if the model is certain of its decision. The overall accuracy could be improved by using ensembles of embeddings, where the outputs are ranked by their (un)certainty.